# Percolation on graphs of polynomial growth is local: analyticity, supercritical sharpness, isoperimetry

### v1

@@ Abstract
This includes uniform versions of supercritical sharpness as well as the Kesten–Zhang bound on the probability of observing a large finite cluster: the constants involved can be chosen to depend only on $(\mathscr{G},p)$. We also prove that $\theta_{\mathscr{H}}$ is an analytic function of $p$ in the whole supercritical regime and that, for a suitable $\varepsilon=\varepsilon(\mathscr{G},p)>0$, the analytic extension of $\theta_{\mathscr{H}}$ to the $\varepsilon$-neighbourhood of $p$ in $\mathbb{C}$ is, uniformly, well approximated by the analytic extension of $\theta_{\mathscr{G}}$. The proof relies on new results on the connectivity of minimal cutsets; in particular, we answer a question asked by Babson and Benjamini in 1999. We further discuss connections with the conjecture of non-percolation at criticality.
@@ 1 Introduction
Likewise, a sharp upper bound on the probability that the cluster of the origin is finite but has cardinality at least $n$ was obtained in [CMT24, Theorem 1.2] and we prove that these estimates hold uniformly on a neighbourhood of $\mathscr{G}$ in $\mathfrak{G}$; see Theorem 4.2 in Section 4.2.
Our techniques go beyond such inequalities and can be used to prove convergence of diverse percolation-related quantities, including the function $\theta$. It is classical in statistical mechanics to define a phase transition to be a parameter where an observable describing the macroscopic behaviour of the system fails to be analytic. For percolation theory, the standard choice is $\theta$, and indeed $\theta$ witnesses the phase transition of the model at $p_{c}$. It was recently proved in [GP23] that for the hypercubic lattice $\mathbb{Z}^{d}$, its cluster-density function $\theta_{\mathscr{G}}$ is analytic on $(p_{c},1]$, thereby answering an old question of Kesten [Kes81]. Our first result in this direction extends this work to all graphs in $\mathfrak{G}$, thus verifying that $\theta$ does not witness any other phase transitions.
@@
Continuity of the map $p_{c}:\mathfrak{G}\to[0,1]$ was originally derived in [CMT23] by using the main results of [CMT24]. Our proof of Theorem 1.4 provides an alternative way to perform the derivation. Indeed, pointwise convergence of $\theta_{\mathscr{G}_{n}}$ to $\theta_{\mathscr{G}_{\infty}}$ on $(p_{c}(\mathscr{G}_{\infty}),1]$ implies upper semicontinuity of $p_{c}:\mathfrak{G}\to[0,1]$. Lower semicontinuity is an older result known to hold for all transitive graphs; see [Pet, Section 14.2] or [DT16, Section 1.2].
@@
Let $\mathscr{G}$ be a graph. Let $u,v\in V(\mathscr{G})$ and $\Pi\subset V(\mathscr{G})$. We say that $\Pi$ is a cutset between $u$ and $v$ if any path from $u$ to $v$ has to pass through at least one vertex of $\Pi$. We say that $\Pi$ is a minimal cutset between $u$ and $v$ if it is minimal for inclusion among all cutsets between $u$ and $v$. Likewise, by taking $\Pi\subset E(\mathscr{G})$, we can define the notions of bond-cutsets and minimal bond-cutsets. For $k\geqslant 1$ and $\Pi\subset V(\mathscr{G})$, we say that $\Pi$ is $k$-connected if it is connected as a subset of $\left(V,E_{k}\right)$, where
@@
It follows from works of Babson–Benjamini and Timár [BB99, Tim07] that for every $\mathscr{G}\in\mathfrak{G}$, there is a finite constant $c$ such that for all vertices $u$ and $v$, all minimal cutsets between $u$ and $v$ are $c$-connected; see [CMT24, Lemma 2.1]. Our next theorem is a local version of this result for finite minimal cutsets.
@@
For every $\mathscr{G}\in\mathfrak{G}$, there is some $c<\infty$ such that the set of all $\mathscr{H}\in\mathfrak{G}$ satisfying the following condition is a neighbourhood of $\mathscr{G}$: for any two vertices $u,v\in V(\mathscr{H})$, every finite minimal cutset from $u$ to $v$ is $c$-connected. $\downarrow$ proof
@@
Let us now come back to percolation and end this introduction by focusing once again on the properties of the percolation function $\theta$. One can formulate a family of interesting problems in percolation theory by investigating the regularity of $\theta$. Depending on the meaning one gives to the word “regularity”, and depending on the graphs one considers, this can give rise to diverse mathematical problems. When one fixes a graph $\mathscr{G}$, the continuity of $\theta$ at $p_{c}(\mathscr{G})$ is equivalent to the absence of percolation at criticality, i.e. whether $\theta(p_{c})=0$, a famous open problem which has been studied extensively in the case of the hypercubic lattice $\mathbb{Z}^{d}$. Continuity at $p_{c}$ is known in dimension $d=2$ [Kes80] and in dimensions $d\geqslant 11$ [HS90, FvdH17], but remains open in all intermediate dimensions, with the case of dimension $d=3$ being notoriously difficult. See [BLPS99, Tim06, PPS06, DCST16, Hut16, HH21] for related works beyond the context of $\mathbb{Z}^{d}$. On the other hand, showing continuity of $\theta$ on $0,1]\setminus\{p_{c}\}$ is much more tractable; see [Sch99] for a proof holding for all transitive graphs.
@@
We are now able to state the following proposition, which is in the same spirit as [DCST16, Proposition 3]. Equivalence of the first two items is classical but we restate it for emphasis.
@@ Acknowledgements
We thank Vincent Tassion for an insightful conversation that inspired the content of Section 6. We are also grateful to Tom Hutchcroft, Ádám Timár, and Matthew Tointon for helpful discussions. CP was supported by an EPSRC New Investigator Award (UKRI1019).
@@ 2 Minimal cutsets are coarsely connected
We denote by $\mathbf{C}(\mathscr{G})$ the infimum of all $k\geqslant 1$ such that, for every $u,v\in V$, every minimal cutset between $u$ and $v$ is $k$-connected. If no such $k$ exists, we set $\mathbf{C}(\mathscr{G})=\infty$. This quantity has been studied in [BB99, Tim07, Tim13], and it will play an important role in this paper. We will also need a variant of this quantity. We denote by $\mathbf{C}^{\star}(\mathscr{G})$ the infimum of all $k\geqslant 1$ such that, for every $u,v\in V$, every finite minimal cutset between $u$ and $v$ is $k$-connected. Again, if no such $k$ exists, we set $\mathbf{C}^{\star}(\mathscr{G})=\infty$.
@@ 2.2 Cutsets and bond-cutsets
Let $\Pi$ be a finite cutset between two vertices $u$ and $v$. We will show that $\Pi$ is $(\mathbf{C}^{\star}_{E}(\mathscr{G})+2)$-connected. If $\Pi=\{u\}$ or $\Pi=\{v\}$, then there is nothing to prove, so let us assume that $\Pi\neq\{u\},\{v\}$.
Let $D$ be the connected component of $u$ in the complement of $\Pi$. Note that $\partial_{E}D$ is a finite bond-cutset that separates $u$ and $v$. Furthermore, the minimality of $\Pi$ implies that $\partial_{E}D$ is a minimal bond-cutset, and that each vertex in $\Pi$ is an endpoint of some edge in $\partial_{E}D$. Hence $\partial_{E}D$ is $\mathbf{C}^{\star}_{E}(\mathscr{G})$-connected, and by the triangle inequality, $\Pi$ is $(\mathbf{C}^{\star}_{E}(\mathscr{G})+2)$-connected, as desired. ∎
@@ 2.3 Cutconnectivity and fibrations
Let $\Pi$ be a finite minimal bond-cutset between two vertices $u$ and $v$ in $\mathscr{S}$. We wish to prove that $\Pi$ is $\mathbf{C}_{E}(\mathscr{L})$-connected. As $\Pi$ is a finite minimal bond-cutset and because $\mathscr{S}$ is one-ended, the graph $\mathscr{S}\setminus\Pi$ must consist of exactly two connected components: a finite one, with vertex-set $F$, and an infinite one, with vertex-set $K$. This is where we use the fact that we work with bond-cutsets instead of cutsets. Notice that $\partial_{E}F=\Pi$ and $\partial_{E}K=\Pi$. As every edge of $\Pi$ has an endpoint in $F$ and the other in $K$, the graph structure induced by $\mathscr{S}$ and by $\mathscr{S}\setminus\Pi$ on $F$ are the same, and so is the case for $K$.
@@ 2.4 Cutconnectivity and quasi-isometries
Let $A$ and $B$ be two constants. Let also $\mathscr{G}$ and $\mathscr{H}$ be two graphs as in the statement. Consider maps $\varphi:V(\mathscr{G})\to V(\mathscr{H})$ and $\psi:V(\mathscr{H})\to V(\mathscr{G})$ that are $A$-quasi-isometries and $A$-quasi-inverse of one another. Pick ${C^{\prime}}\geqslant A^{2}+2A$, ${C^{\prime\prime}}\geqslant A(A+2{C^{\prime}}+B+3)$ and $C=2{C^{\prime\prime}}+1$. Let $\Pi$ be a finite minimal cutset between two vertices $x$ and $y$ in $\mathscr{H}$. Our goal is to prove that $\Pi$ is $C$-connected.
Proceeding by contradiction, we assume that $\Pi$ is not $C$-connected. Let us denote by $K_{1}$ a $C$-connected component of $\Pi$, and by $K_{2}$ the union of all other $C$-connected components of $\Pi$. By construction, and since we assume $\Pi$ not to be $C$-connected, $(K_{1},K_{2})$ is a partition of $\Pi$ into two non-empty subsets, with $K_{1}$ being $C$-connected.
@@
Let us now prove that there is a path from $x^{\prime}$ to $y^{\prime}$ that avoids $K_{2}({C^{\prime\prime}})$. By definition of $x^{\prime}$, we can pick a neighbour $x^{\prime\prime}$ of $x^{\prime}$ that belongs to $K_{1}({C^{\prime\prime}})$. Likewise, pick a neighbour $y^{\prime\prime}$ of $y^{\prime}$ that belongs to $K_{1}({C^{\prime\prime}})$. Since $K_{1}$ is $C$-connected and $C\leqslant 2{C^{\prime\prime}}+1$, the set $K_{1}({C^{\prime\prime}})$ is connected. Moreover, as it is disjoint from $K_{2}({C^{\prime\prime}})$, we can pick a path from $x^{\prime\prime}$ to $y^{\prime\prime}$ that avoids $K_{2}({C^{\prime\prime}})$. Appending $x^{\prime}$ at the beginning and $y^{\prime}$ at the end of this path yields a path from $x^{\prime}$ to $y^{\prime}$ that avoids $K_{2}({C^{\prime\prime}})$.
@@
We can now conclude as follows. The set $\hat{\Pi}=\hat{K_{1}}\cup\hat{K_{2}}$ is a cutset between $u$ and $v$ and none of $\hat{K_{1}}$ and $\hat{K_{2}}$ is a cutset between $u$ and $v$. As a consequence, there is a minimal cutset contained in $\hat{\Pi}$ that intersects both $\hat{K_{1}}$ and $\hat{K_{2}}$. Therefore, we can pick $w_{1}\in\hat{K_{1}}$ and $w_{2}\in\hat{K_{2}}$ such that $d_{\mathscr{G}}(w_{1},w_{2})\leqslant B$. By definition of $\hat{K_{1}}$ and $\hat{K_{2}}$, we can further pick $z^{\prime}_{1}\in K_{1}$ and $z^{\prime}_{2}\in K_{2}$ such that $d_{\mathscr{G}}(w_{1},\psi(z^{\prime}_{1}))\leqslant{C^{\prime}}$ and $d_{\mathscr{G}}(w_{2},\psi(z^{\prime}_{2}))\leqslant{C^{\prime}}$. Using the triangle inequality, we get $d_{\mathscr{G}}(\psi(z^{\prime}_{1}),\psi(z^{\prime}_{2}))\leqslant 2{C^{\prime}}+B$. Applying $\varphi$ and recalling that $\varphi\circ\psi$ is $A$-close to the identity, we get $d_{\mathscr{H}}(z^{\prime}_{1},z^{\prime}_{2})\leqslant(2{C^{\prime}}+B)A+3A\leqslant C$. This contradicts the fact that $K_{2}$ is disjoint from the $C$-connected component $K_{1}$ of $\Pi$, thus concluding the proof. ∎
Remark 2.15.
In our proof of Proposition 2.12, the constant $C^{\prime}$} is taken large enough and $C^{\prime\prime}$} large enough depending on $C^{\prime}$}. However, regarding the value of $C$, we need it to satisfy inequalities in two opposite directions, namely $C\leqslant 2{C^{\prime\prime}}+1$ and $C\geqslant 2{C^{\prime\prime}}+1$. Fortunately, one may pick $C=2{C^{\prime\prime}}+1$ so that both conditions are met.
@@
Let us fix a constant $t$ such that the conclusion of Theorem 2.2 holds. In particular, $\partial\mathcal{C}_{o}(N)$ is $t$-connected for every $N\geqslant 1$. Below, we will take $N$ to be sufficiently large, and in particular choose $N\geqslant t$.
@@
Let $N\geqslant t$, and let $\omega$ be a configuration such that $20N<\mathrm{diam}(\mathcal{C}_{o})<\infty$. Define $\mathcal{I}$ to be the $t$-connected component of $\partial\mathcal{C}_{o}(N)$ within the set of $N$-bad vertices in $\omega$. Define $\overline{\mathcal{I}}$ to be the set of vertices $x\in V$ such that $d_{\mathscr{G}}(x,\mathcal{I})\leqslant 10N$.
@@
Now we connect $u$ to $v$ by an open path as follows. Since $\mathbf{C}^{\star}(\mathscr{G})\leqslant t$, we know that $\partial A$ is $t$-connected, which means that we can find a sequence of vertices $v_{1}=w^{\prime},\ldots,v_{\ell}=v$ in $\partial A$ such that $d_{\mathscr{G}}(v_{i},v_{i+1})\leqslant t$ for every $i<\ell$. As all vertices $v_{1},\ldots,v_{\ell}$ are $N$-good, $B_{N}(v_{i})$ is connected to $S_{10N}(v_{i})$ by an open for every $i\leqslant\ell$. Furthermore, $w$ is connected to $S_{5N}(w^{\prime})$ by an open path due to the fact that $\mathrm{diam}(\mathcal{C}_{w})>20N$, and by (4), the vertex $v=v_{\ell}$ is connected to $S_{5N}(v_{\ell})$ by an open path. Using that $N\geqslant t$ and that all the events $U(2N,5N,v_{i})$ happen, we deduce that we can connect $w$ to $v_{\ell}=v$ by an open path. Since $w\in\mathcal{C}_{o}$, we can thus connect $u$ to $v$ by an open path, which proves the claim and completes the proof. ∎
@@
Let $\mathscr{I}\subset V$ be a finite set of vertices. We say that $\mathscr{I}$ is an interface if it is $t$-connected and either $o\in\mathscr{I}$ or $o$ lies in a finite connected components of $\mathscr{G}\setminus\mathscr{I}$. Let $\overline{\mathscr{I}}$ denote the set of vertices $x$ such that $d_{\mathscr{G}}(x,\mathscr{I})\leqslant 5N$. We say that an interface $\mathscr{I}$ occurs in $\omega$ at scale $N$ if
@@
Since interfaces are $t$-connected and because of items 1 and 2 in Definition 3.4, occurring interfaces are in particular $t$-connected components of the set of bad vertices. The desired result follows. ∎
Let us write $\mathcal{C}_{x}^{\mathrm{bad}}(N)$ for the $t$-connected component of $x$ in $\omega$ within the set of $N$-bad vertices, with the convention that $\mathcal{C}_{x}^{\mathrm{bad}}(N)=\varnothing$ if $x$ is $N$-good. We now show that for every $p>p_{c}$, for every $N$ large enough (depending on $p$), $\mathcal{C}_{x}^{\mathrm{bad}}(N)$ is almost surely finite, which will allow us to deduce that $\overline{\mathcal{I}}$ is almost surely finite. The proof is similar to that of [CMT24, Lemma 10.1]. We include a proof for completeness and to justify the uniformity in $p$ in the statement. The latter is crucial for obtaining a suitable expression for $\theta$ that holds in an open neighbourhood.

### v2

# arXiv:2511.01851v2[math.PR] 01 Mar 2026

@@ Abstract
This includes uniform versions of supercritical sharpness as well as the Kesten–Zhang bound on the probability of observing a large finite cluster and the exponential cluster repulsion bound of Pete: the constants involved can be chosen to depend only on $(\mathscr{G},p)$. We also prove that $\theta_{\mathscr{H}}$ is an analytic function of $p$ in the whole supercritical regime and that, for a suitable $\varepsilon=\varepsilon(\mathscr{G},p)>0$, the analytic extension of $\theta_{\mathscr{H}}$ to the $\varepsilon$-neighbourhood of $p$ in $\mathbb{C}$ is, uniformly, well approximated by the analytic extension of $\theta_{\mathscr{G}}$. The proof relies on new results on the connectivity of minimal cutsets; in particular, we answer a question asked by Babson and Benjamini in 1999. We further discuss connections with the conjecture of non-percolation at criticality.
@@ 1 Introduction
Likewise, a sharp upper bound on the probability that the cluster of the origin is finite but has cardinality at least $n$ was obtained in [CMT24, Theorem 1.2]. We prove that these estimates hold uniformly on a neighbourhood of $\mathscr{G}$ in $\mathfrak{G}$; see Theorem 4.2 in Section 4.2.
Beyond the cluster size distribution, of particular interest is the isoperimetry of $\mathcal{C}_{o}$. In Theorem 4.5, we prove that when $\mathcal{C}_{o}$ is finite, the number of edges that touch both $\mathcal{C}_{o}$ and the infinite cluster has an exponential tail, uniformly on a neighbourhood of $\mathscr{G}$ in $\mathfrak{G}$. This extends an earlier result of Pete [PET08] from the context of $\mathbb{Z}^{d}$ to that of $\mathfrak{G}$. Using this exponential cluster repulsion result together with the stretched exponential decay of the cluster size, and following the argument of Pete [PET08], we show that the isoperimetric profile of the infinite cluster roughly coincides with that of the original lattice. As established in [PET08] in the case of $\mathbb{Z}^{d}$, this has important implications, such as that simple random walk on the largest cluster of a finite box $n,n]^{d}$ has, with high probability, $L^{\infty}$-mixing time $\Theta(n^{2})$, and that the return probability $p_{n}(o,o)$ on the infinite cluster almost surely decays as $p_{n}(o,o)=O(n^{-d/2})$. We expect all these random walk results to carry over to the context of $\mathfrak{G}$, but verifying this is beyond the scope of the present paper.
Our techniques go beyond such results and can be used to prove convergence of diverse percolation-related quantities, including the function $\theta$. It is classical in statistical mechanics to define a phase transition to be a parameter where an observable describing the macroscopic behaviour of the system fails to be analytic. For percolation theory, the standard choice is $\theta$, and indeed $\theta$ witnesses the phase transition of the model at $p_{c}$. It was recently proved in [GP23] that for the hypercubic lattice $\mathbb{Z}^{d}$, its cluster-density function $\theta_{\mathscr{G}}$ is analytic on $(p_{c},1]$, thereby answering an old question of Kesten [KES81]. Our first result in this direction extends this work to all graphs in $\mathfrak{G}$, thus verifying that $\theta$ does not witness any other phase transitions.
@@
Continuity of the map $p_{c}:\mathfrak{G}\to[0,1]$ was originally derived in [CMT23] by using the main results of [CMT24]. Our proof of Theorem 1.4 provides an alternative way to perform the derivation. Indeed, pointwise convergence of $\theta_{\mathscr{G}_{n}}$ to $\theta_{\mathscr{G}_{\infty}}$ on $(p_{c}(\mathscr{G}_{\infty}),1]$ implies upper semicontinuity of $p_{c}:\mathfrak{G}\to[0,1]$. Lower semicontinuity is an older result known to hold for all transitive graphs; see [36, Section 14.2] or [DT16, Section 1.2].
@@
Let $\mathscr{G}$ be a graph. Let $u,v\in V(\mathscr{G})$ and $\Pi\subset V(\mathscr{G})$. We say that $\Pi$ is a cutset between $u$ and $v$ if any path from $u$ to $v$ has to pass through at least one vertex of $\Pi$. We say that $\Pi$ is a minimal cutset between $u$ and $v$ if it is minimal for inclusion among all cutsets between $u$ and $v$. Likewise, by taking $\Pi\subset E(\mathscr{G})$, we can define the notions of bond-cutsets and minimal bond-cutsets. For $k\geqslant 1$ and $\Pi\subset V(\mathscr{G})$, we say that $\Pi$ is $k$-coarse connected if it is connected as a subset of $\left(V,E_{k}\right)$, where
@@
It follows from works of Babson–Benjamini and Timár [BB99, TIM07] that for every $\mathscr{G}\in\mathfrak{G}$, there is a finite constant $c$ such that for all vertices $u$ and $v$, all minimal cutsets between $u$ and $v$ are $c$-coarse connected; see [CMT24, Lemma 2.1]. Our next theorem is a local version of this result for finite minimal cutsets.
@@
For every $\mathscr{G}\in\mathfrak{G}$, there is some $c<\infty$ such that the set of all $\mathscr{H}\in\mathfrak{G}$ satisfying the following condition is a neighbourhood of $\mathscr{G}$: for any two vertices $u,v\in V(\mathscr{H})$, every finite minimal cutset from $u$ to $v$ is $c$-coarse connected. $\downarrow$ proof
@@
Remark 1.9.
It is a standard fact that, on connected graphs of bounded degree, coarse connectivity of cutsets implies an exponential upper bound on the number of cutsets of a given size that separate a fixed origin from infinity. Our Theorem 1.8 implies that this exponential upper bound is local for graphs in $\mathfrak{G}$. The latter can also be deduced from the results of [EST24], namely from the fact that $\theta_{\mathscr{G}}(1-\varepsilon)\geqslant\varepsilon$ for some uniform $\varepsilon>0$, which follows from the proof of the so-called gap at $1$ for $p_{c}$, and their Theorem $1$. We remark that when combined with the results of [CMT24], this is sufficient to prove Theorem 1.2. However, for the proof of the analyticity of $\theta$, it is crucial that the coarse connectivity of cutsets is itself local.
Let us now come back to percolation and end this introduction by focusing once again on the properties of the percolation function $\theta$. One can formulate a family of interesting problems in percolation theory by investigating the regularity of $\theta$. Depending on the meaning one gives to the word “regularity”, and depending on the graphs one considers, this can give rise to diverse mathematical problems. When one fixes a graph $\mathscr{G}$, the continuity of $\theta$ at $p_{c}(\mathscr{G})$ is equivalent to the absence of percolation at criticality, i.e. whether $\theta(p_{c})=0$, a famous open problem which has been studied extensively in the case of the hypercubic lattice $\mathbb{Z}^{d}$. Continuity at $p_{c}$ is known in dimension $d=2$ [KES80] and in dimensions $d\geqslant 11$ [HS90, Fv17], but remains open in all intermediate dimensions, with the case of dimension $d=3$ being notoriously difficult. See [BLP+99, TIM06, PPS06, DST16, HUT16, HH21] for related works beyond the context of $\mathbb{Z}^{d}$. On the other hand, showing continuity of $\theta$ on $0,1]\setminus\{p_{c}\}$ is much more tractable; see [SCH99] for a proof holding for all transitive graphs.
@@
We are now able to state the following proposition, which is in the same spirit as [DST16, Proposition 3]. Equivalence of the first two items is classical but we restate it for emphasis.
@@ Acknowledgements
We thank Vincent Tassion for an insightful conversation that inspired the content of Section 6. We are grateful to Gábor Pete for suggesting to include a proof of local forms of exponential cluster repulsion and anchored-isoperimetric estimates. We thank Franco Severo for pointing out that [EST24] gives a local bound on the number of minimal cutsets. At last, we are grateful to Nathan Deloire, Alexander Glazman, Tom Hutchcroft, Ádám Timár, and Matthew Tointon for helpful discussions. CP was supported by an EPSRC New Investigator Award (UKRI1019).
@@ 2 Minimal cutsets are coarsely connected
We denote by $\mathbf{C}(\mathscr{G})$ the infimum of all $k\geqslant 1$ such that, for every $u,v\in V$, every minimal cutset between $u$ and $v$ is $k$-coarse connected. If no such $k$ exists, we set $\mathbf{C}(\mathscr{G})=\infty$. This quantity has been studied in [BB99, TIM07, TIM13], and it will play an important role in this paper. We will also need a variant of this quantity. We denote by $\mathbf{C}^{\star}(\mathscr{G})$ the infimum of all $k\geqslant 1$ such that, for every $u,v\in V$, every finite minimal cutset between $u$ and $v$ is $k$-coarse connected. Again, if no such $k$ exists, we set $\mathbf{C}^{\star}(\mathscr{G})=\infty$.
@@ 2.2 Cutsets and bond-cutsets
Let $\Pi$ be a finite cutset between two vertices $u$ and $v$. We will show that $\Pi$ is $(\mathbf{C}^{\star}_{E}(\mathscr{G})+2)$-coarse connected. If $\Pi=\{u\}$ or $\Pi=\{v\}$, then there is nothing to prove, so let us assume that $\Pi\neq\{u\},\{v\}$.
Let $D$ be the connected component of $u$ in the complement of $\Pi$. Note that $\partial_{E}D$ is a finite bond-cutset that separates $u$ and $v$. Furthermore, the minimality of $\Pi$ implies that $\partial_{E}D$ is a minimal bond-cutset, and that each vertex in $\Pi$ is an endpoint of some edge in $\partial_{E}D$. Hence $\partial_{E}D$ is $\mathbf{C}^{\star}_{E}(\mathscr{G})$-coarse connected, and by the triangle inequality, $\Pi$ is $(\mathbf{C}^{\star}_{E}(\mathscr{G})+2)$-coarse connected, as desired. ∎
@@ 2.3 Cutconnectivity and fibrations
Let $\Pi$ be a finite minimal bond-cutset between two vertices $u$ and $v$ in $\mathscr{S}$. We wish to prove that $\Pi$ is $\mathbf{C}_{E}(\mathscr{L})$-coarse connected. As $\Pi$ is a finite minimal bond-cutset and because $\mathscr{S}$ is one-ended, the graph $\mathscr{S}\setminus\Pi$ must consist of exactly two connected components: a finite one, with vertex-set $F$, and an infinite one, with vertex-set $K$. This is where we use the fact that we work with bond-cutsets instead of cutsets. Notice that $\partial_{E}F=\Pi$ and $\partial_{E}K=\Pi$. As every edge of $\Pi$ has an endpoint in $F$ and the other in $K$, the graph structure induced by $\mathscr{S}$ and by $\mathscr{S}\setminus\Pi$ on $F$ are the same, and so is the case for $K$.
@@ 2.4 Cutconnectivity and quasi-isometries
Let $A$ and $B$ be two constants. Let also $\mathscr{G}$ and $\mathscr{H}$ be two graphs as in the statement. Consider maps $\varphi:V(\mathscr{G})\to V(\mathscr{H})$ and $\psi:V(\mathscr{H})\to V(\mathscr{G})$ that are $A$-quasi-isometries and $A$-quasi-inverse of one another. Pick ${C^{\prime}}\geqslant A^{2}+2A$, ${C^{\prime\prime}}\geqslant A(A+2{C^{\prime}}+B+3)$ and $C=2{C^{\prime\prime}}+1$. Let $\Pi$ be a finite minimal cutset between two vertices $x$ and $y$ in $\mathscr{H}$. Our goal is to prove that $\Pi$ is $C$-coarse connected.
Proceeding by contradiction, we assume that $\Pi$ is not $C$-coarse connected. Let us denote by $K_{1}$ a $C$-coarse connected component of $\Pi$, and by $K_{2}$ the union of all other $C$-coarse connected components of $\Pi$. By construction, and since we assume $\Pi$ not to be $C$-coarse connected, $(K_{1},K_{2})$ is a partition of $\Pi$ into two non-empty subsets, with $K_{1}$ being $C$-coarse connected.
@@
Let us now prove that there is a path from $x^{\prime}$ to $y^{\prime}$ that avoids $K_{2}({C^{\prime\prime}})$. By definition of $x^{\prime}$, we can pick a neighbour $x^{\prime\prime}$ of $x^{\prime}$ that belongs to $K_{1}({C^{\prime\prime}})$. Likewise, pick a neighbour $y^{\prime\prime}$ of $y^{\prime}$ that belongs to $K_{1}({C^{\prime\prime}})$. Since $K_{1}$ is $C$-coarse connected and $C\leqslant 2{C^{\prime\prime}}+1$, the set $K_{1}({C^{\prime\prime}})$ is connected. Moreover, as it is disjoint from $K_{2}({C^{\prime\prime}})$, we can pick a path from $x^{\prime\prime}$ to $y^{\prime\prime}$ that avoids $K_{2}({C^{\prime\prime}})$. Appending $x^{\prime}$ at the beginning and $y^{\prime}$ at the end of this path yields a path from $x^{\prime}$ to $y^{\prime}$ that avoids $K_{2}({C^{\prime\prime}})$.
@@
We can now conclude as follows. The set $\hat{\Pi}=\hat{K_{1}}\cup\hat{K_{2}}$ is a cutset between $u$ and $v$ and none of $\hat{K_{1}}$ and $\hat{K_{2}}$ is a cutset between $u$ and $v$. As a consequence, there is a minimal cutset contained in $\hat{\Pi}$ that intersects both $\hat{K_{1}}$ and $\hat{K_{2}}$. Therefore, we can pick $w_{1}\in\hat{K_{1}}$ and $w_{2}\in\hat{K_{2}}$ such that $d_{\mathscr{G}}(w_{1},w_{2})\leqslant B$. By definition of $\hat{K_{1}}$ and $\hat{K_{2}}$, we can further pick $z^{\prime}_{1}\in K_{1}$ and $z^{\prime}_{2}\in K_{2}$ such that $d_{\mathscr{G}}(w_{1},\psi(z^{\prime}_{1}))\leqslant{C^{\prime}}$ and $d_{\mathscr{G}}(w_{2},\psi(z^{\prime}_{2}))\leqslant{C^{\prime}}$. Using the triangle inequality, we get $d_{\mathscr{G}}(\psi(z^{\prime}_{1}),\psi(z^{\prime}_{2}))\leqslant 2{C^{\prime}}+B$. Applying $\varphi$ and recalling that $\varphi\circ\psi$ is $A$-close to the identity, we get $d_{\mathscr{H}}(z^{\prime}_{1},z^{\prime}_{2})\leqslant(2{C^{\prime}}+B)A+3A\leqslant C$. This contradicts the fact that $K_{2}$ is disjoint from the $C$-coarse connected component $K_{1}$ of $\Pi$, thus concluding the proof. ∎
@@
Let us fix a constant $t$ such that the conclusion of Theorem 2.2 holds. In particular, $\partial\mathcal{C}_{o}(N)$ is $t$-coarse connected for every $N\geqslant 1$. Below, we will take $N$ to be sufficiently large, and in particular choose $N\geqslant t$.
@@
Let $N\geqslant t$, and let $\omega$ be a configuration such that $20N<\mathrm{diam}(\mathcal{C}_{o})<\infty$. Define $\mathcal{I}$ to be the $t$-coarse connected component of $\partial\mathcal{C}_{o}(N)$ within the set of $N$-bad vertices in $\omega$. Define $\overline{\mathcal{I}}$ to be the set of vertices $x\in V$ such that $d_{\mathscr{G}}(x,\mathcal{I})\leqslant 10N$.
@@
Now we connect $u$ to $v$ by an open path as follows. Since $\mathbf{C}^{\star}(\mathscr{G})\leqslant t$, we know that $\partial A$ is $t$-coarse connected, which means that we can find a sequence of vertices $v_{1}=w^{\prime},\ldots,v_{\ell}=v$ in $\partial A$ such that $d_{\mathscr{G}}(v_{i},v_{i+1})\leqslant t$ for every $i<\ell$. As all vertices $v_{1},\ldots,v_{\ell}$ are $N$-good, $B_{N}(v_{i})$ is connected to $S_{10N}(v_{i})$ by an open for every $i\leqslant\ell$. Furthermore, $w$ is connected to $S_{5N}(w^{\prime})$ by an open path due to the fact that $\mathrm{diam}(\mathcal{C}_{w})>20N$, and by (4), the vertex $v=v_{\ell}$ is connected to $S_{5N}(v_{\ell})$ by an open path. Using that $N\geqslant t$ and that all the events $U(2N,5N,v_{i})$ happen, we deduce that we can connect $w$ to $v_{\ell}=v$ by an open path. Since $w\in\mathcal{C}_{o}$, we can thus connect $u$ to $v$ by an open path, which proves the claim and completes the proof. ∎
@@
Let $\mathscr{I}\subset V$ be a finite set of vertices. We say that $\mathscr{I}$ is an interface if it is $t$-coarse connected and either $o\in\mathscr{I}$ or $o$ lies in a finite connected components of $\mathscr{G}\setminus\mathscr{I}$. Let $\overline{\mathscr{I}}$ denote the set of vertices $x$ such that $d_{\mathscr{G}}(x,\mathscr{I})\leqslant 10N$. We say that an interface $\mathscr{I}$ occurs in $\omega$ at scale $N$ if
@@
Since interfaces are $t$-coarse connected and because of items 1 and 2 in Definition 3.4, occurring interfaces are in particular $t$-coarse connected components of the set of bad vertices. The desired result follows. ∎
Let us write $\mathcal{C}_{x}^{\mathrm{bad}}(N)$ for the $t$-coarse connected component of $x$ in $\omega$ within the set of $N$-bad vertices, with the convention that $\mathcal{C}_{x}^{\mathrm{bad}}(N)=\varnothing$ if $x$ is $N$-good. We now show that for every $p>p_{c}$, for every $N$ large enough (depending on $p$), $\mathcal{C}_{x}^{\mathrm{bad}}(N)$ is almost surely finite, which will allow us to deduce that $\overline{\mathcal{I}}$ is almost surely finite. The proof is similar to that of [CMT24, Lemma 10.1]. We include a proof for completeness and to justify the uniformity in $p$ in the statement. The latter is crucial for obtaining a suitable expression for $\theta$ that holds in an open neighbourhood.